# Regulating radiology AI medical devices that evolve in their lifecycle

### v2

@@ Abstract
Over time, the distribution of medical image data drifts due to multiple factors, including shifts in patient demographics, acquisition devices, and disease manifestation. While human radiologists can extrapolate their knowledge to such changes, AI systems cannot. In fact, deep learning models are highly susceptible to even slight variations in image characteristics. Therefore, manufacturers must update their models with new data to ensure that they remain safe and effective. Until recently, conducting such model updates in the USA and European Union meant applying for re-approval. Given the time and monetary costs associated with these processes, updates were infrequent, and obsolete systems continued functioning for too long. During 2024, several developments in the regulatory frameworks of these regions have taken place that promise to streamline the process of rolling out model updates safely: The European Artificial Intelligence Act came into effect last August, and the Food and Drug Administration (FDA) released the final marketing submission recommendations for a Predetermined Change Control Plan (PCCP) in December. We give an overview of the requirements and objectives of recent regulatory efforts and summarize the building blocks needed for successfully deploying dynamic systems. At the center of these pieces of regulation – and as prerequisites for manufacturers to conduct model updates without re-approval – are the need to describe the data collection and re-training processes and to establish real-world quality monitoring mechanisms.
@@ 1 Introduction
Imagine a use case where an AI-supported system could be helpful in the clinical routine, such as detection of pulmonary embolism in chest CT (Colak et al., 2021). The process from design to deployment requires a complex vertical integration from collecting data to embedding the workflow in clinical practice (Fuchs et al., 2023). The product – or Software as a Medical Device (SaMD) (IMDRF, 2013) –would typically go through five stages, from design to deployment:
@@
This process seems reasonable, and the product will initially perform well as long as the evaluation has been sufficiently thorough and suitable for the clinical problem. What happens, however, after one year? How about five?
@@ 3 Building an infrastructure for continuous AI
Figure 4: From left to right: a radiologist receives a chest CT scan and produces an initial diagnosis for pulmonary embolism. In parallel, a UNet extracts a segmentation which, if deemed reliable by the QA system, is delivered to the radiologist alongside saliency maps. With the new input, the doctor rectifies the report by adding small emboli that were initially not detected; and later corrects the segmentation mask for a portion of the volume. The image and corrected segmentation can then be used to improve model performance as an additional data sample.
@@
Of course, maintaining an in-house dataset for benchmarking and manual quality monitoring are tasks that should be carried out for locked just as for lifelong products. In addition, active learning methods can produce high-quality annotations with minimal user input (Gotkowski et al., 2022). Figure 4 illustrates how active learning and interpretability techniques can be used to improve user experience and decrease the workload of radiologists.
@@ 4 Conclusions and Outlook
Until recently, manufacturers of AI medical software were reluctant to develop products that continued learning post-deployment due to the regulatory overhead involved with obtaining re-certification. Instead, they would stand on large and heterogeneous datasets hoping that models would stay robust for as long as possible. But the risk of progressive performance deterioration of locked AI cannot be overlooked.

### v3

@@ Abstract
Over time, the distribution of medical image data drifts due to factors such as shifts in patient demographics, acquisition devices, and disease manifestations. While human radiologists can adjust their expertise to accommodate such variations, deep learning models cannot. In fact, such models are highly susceptible to even slight variations in image characteristics. Consequently, manufacturers must conduct regular updates to ensure that they remain safe and effective. Performing such updates in the United States and European Union required, until recently, obtaining re-approval. Given the time and financial burdens associated with these processes, updates were infrequent, and obsolete systems remained in operation for too long. During 2024, several regulatory developments promised to streamline the safe rollout of model updates: The European Artificial Intelligence Act came into effect last August, and the Food and Drug Administration (FDA) issued final marketing submission recommendations for a Predetermined Change Control Plan (PCCP) in December. We provide an overview of these developments and outline the key building blocks necessary for successfully deploying dynamic systems. At the heart of these regulations – and as prerequisites for manufacturers to conduct model updates without re-approval – are clear descriptions of data collection and re-training processes, coupled with robust real-world quality monitoring mechanisms.
keywords:
AI medical devices; AI regulation; continual learning
@@ 1 Introduction
Imagine a use case where an AI-supported system could be helpful in the clinical routine, such as detecting pulmonary embolism in chest computer tomography (CT) (Colak et al., 2021). The process from design to deployment requires a complex vertical integration from collecting data to embedding the workflow in clinical practice (Fuchs et al., 2023). The product – or Software as a Medical Device (SaMD) (IMDRF, 2013) –would typically go through the following five stages:
@@
This process seems reasonable, and the product will initially perform well as long as the evaluation was suitable for the clinical problem. What happens, however, after one year? How about five?
@@ 3 Building an infrastructure for continuous AI
Of course, maintaining an in-house dataset for benchmarking and manual quality monitoring are tasks that should be carried out for locked just as for lifelong products. In addition, active learning methods can produce high-quality annotations with minimal user input (Gotkowski et al., 2022).
@@ 4 Conclusions and Outlook
Until recently, manufacturers of AI medical software were reluctant to develop products that continue learning post-deployment due to the regulatory overhead involved with obtaining re-certification. Instead, they would stand on large and heterogeneous datasets hoping that models would stay robust for as long as possible. But the risk of progressive performance deterioration of locked AI cannot be overlooked.